OpenStudy (anonymous):

The octane rating of a gasoline is a measure of the amount of isooctane in the gas. Manufacturers recommend using 93-octane gasoline on returned motors. How much 87-octane gas and 95 octane gas should Yousef mix in order to make 10 gal of 93 octane gas for his returned Ford F-150?

OpenStudy (anonymous):

Hello again, I would like to know how to figure out how to make the equation, then solve

OpenStudy (saifoo.khan):

Hey! We have t make 93 octane and 10 gallons. so our total amount will be 0.93*10.

OpenStudy (saifoo.khan):

Moving along: total gallons will be 10. so 0.95(10-x) Final equation: 0.87x + 0.95(10-x) = 0.93*10

OpenStudy (saifoo.khan):

There's one more method too if this looks hard to you. Let me know if this is hard to understand.

OpenStudy (anonymous):

This method looked a bit more difficult than the latter, do you have an easier method?

OpenStudy (saifoo.khan):

x + y = 10 87x + 95y = 93*10

OpenStudy (anonymous):

I see, I was trying to figure out how you know what goes where?

OpenStudy (saifoo.khan):

Haha. I don't know. I should ask that from my brain! :D

OpenStudy (anonymous):

You are funny! Thank you! I have so much homework to catch up on!I am behind quite a bit, the class is already on Matrices and also sqrt[?]{?}.... I am VERY behind!

OpenStudy (saifoo.khan):

Lol. You're welcome. Oh! Matrices can be boring. Square roots are fun!

OpenStudy (anonymous):

I don't know why but my equation seems wrong... x+y=2 4x+y=5 x=-y+5 -4y+20+y=5 -3y+20=5 -3y=-20+5 -3y=-15 -3y/-3=-15/-3 y=5 x=-5+5 x=0???

OpenStudy (saifoo.khan):

What's the main question?

OpenStudy (anonymous):

How much lowfat milk(1% fat) and how much whole milk(4% fat) should be mixed to make 5 gal of reduced fat milk (2% fat)?

OpenStudy (saifoo.khan):

x + y = 5 0.01x + 0.04y = 0.02*5

OpenStudy (anonymous):

Hmmm... how do you know what to put where?

OpenStudy (saifoo.khan):

It's pretty simple. For equation 1. find the total amount. it will be x+y = 5(total gallons)

OpenStudy (anonymous):

Maybe so simple that I over think and get confused! LOL

OpenStudy (saifoo.khan):

for equation 2. Now we have 3 values left. 1%, 4% and 2% We have to make 2% so: 0.01x + 0.04y = 0.02 *5 *5 because total are 5 gals.

OpenStudy (anonymous):

I see....

OpenStudy (saifoo.khan):

Lol, it happens.

OpenStudy (anonymous):

Where are you located Saifoo? I am so glad you grabbed my questions!

OpenStudy (saifoo.khan):

Lol. I'm located in another continent. Way far from US. I'm from Pakistan.

OpenStudy (anonymous):

Wow... You speak and write English very well! Are you Pakistani or Indian? I am just curious... It is funny how the internet connects people, and we can find common ground and learn from one another....

OpenStudy (anonymous):

Are you in Pakistan now? It is day there I imagine... It is late here in Indiana where I live in the U.S.

OpenStudy (saifoo.khan):

No. I'm a Pakistani. Living in Pakistan currently! :D Hehe. Yup. This world is shorter now. People can connect in seconds. It's 8:14am here. And i should be heading to bed soon. Haven't slept the whole night. :/ Been studying for exams.

OpenStudy (anonymous):

It is very cool to meet someone in another country, again I am glad to know you!

OpenStudy (anonymous):

A 17 year old teaching a 43 year old! WONDERFUL!!! :-D

OpenStudy (saifoo.khan):

Same here. :)

OpenStudy (saifoo.khan):

Haha. Learning anything has no limits.

OpenStudy (anonymous):

Well get some rest, I hope that you are okay with helping me, thank you so much sweetie!

OpenStudy (saifoo.khan):

Yup! Going in 5mins. If my mom sees me awake, i'm dead. D: Yes, sure. Knock me if you need any help. You're welcome ma'am.

OpenStudy (anonymous):

Hey my Facebook is Susan-Jaxie Marie Hanna-Jackson, friend me if you would like! I look forward to learning more from you! Thank you so much again!!!

OpenStudy (saifoo.khan):

just sent you a request. You're welcome. :)